On January 1, 2021, Rupar Retailers purchased $120,000 of Anand Company bonds at a premium of $4,000. The Anand bonds pay 8%Multiple Choice $4,682 $3,688. $3,600. 0 $4,800.
Answer #1
Cash interest received 4800 =120000*8%/2
Interest revenue for first interest payment 3720 =124000*6%/2
Premium amortized for first payment 1080
Carrying value after first payment 122920 =124000-1080
Interest revenue for second interest payment 3688 =122920*6%/2
Option B $3688 is correct